James is as twice as old as his son. in 20 years their combined ages will be 115 years. how old is his son?

Answer by stanbon(75887)

James is as twice as old as his son. in 20 years their combined ages will be 115 years. how old is his son?
----------
Equations:
j = 2*s
j+20 + s+20 = 115
-------
Substitute for "j" and solve for "s":
-------
2s+20 + s+20 = 115
3s + 40 = 115
3s = 75
son's age = 25
----
James's age = 3s = 75
---------------------------
Cheers,
Stan H.
=================